Q1.To find the optimum shape for a cross-section of a beam against bending, beam sections of a square, circle and equilateral triangle with the same cross-sectional area (A) are considered. If the beams have the same length (L), Young’s Modulus (E) and support condition, determine which shape of the beam is the best against bending?
1
Expert's answer
2019-02-26T11:39:58-0500

The deflection of beams in the vertical direction with the length "L", the moment of inertia of cross-section "I" under load "F\/k" is defined by


"\delta=\frac{FL^3}{kEI}."


Thus, since all other properties are equal, the higher "I", the more stress a beam can withstand. So we will just compare "I", higher - better.

For a square cross-section with side "a" it is medium:


"I_{\Box}=\frac{a^4}{12}=\frac{A^2}{12}."


For a circle with radius "r" it is the smallest:


"I_{\bigcirc}=\frac{\pi r^4}{4}=\frac{A^2}{4\pi}=\frac{A^2}{12.57}."


For a triangle with side "b" it is the highest:


"I_{\bigtriangleup}=\frac{b^4}{32\sqrt{3}}=\frac{A^2}{6\sqrt{3}}=\frac{A^2}{10.39}."


Triangles rule.